an object weighing 40 N rests on an inclined plane making an angle of 30 degree with the horizontal. Resolve the force into components parallel to the plane and perpendicular to the plane


1
Expert's answer
2021-07-14T10:16:06-0400
"W_{\text{\textbardbl}}=Wsin\theta=40\ N\cdot sin30^{\circ}=20\ N,""W_{\bot}=Wcos\theta=40\ N\cdot cos30^{\circ}=34.64\ N."